OpenStudy (mera8421):

francis has to prove that a linear function grows by equal differences over equal intervals by using the graph below and the slope-intercept equation, y=mx+b. What is the missing step in the proof? 1. If x1,=a, then y,=ma+b by substituting a into the slope intercept equation. 2. If x2=a+d, then y2=ma+md+bby substituting (a+d) into the slope intercept equation. 3. (triangle)y= y2-y1=__________ by the subtraction property of equality. 4.(triangle)x=x2-x1=(a+d)-a=d by the subtraction of property of equality. 5. using the slope formula, (triangle)y/(triangle)x=y2-y1/x2-x1=md/d=m

OpenStudy (mera8421):

A (ma+md+mb)-(ma+b) B (ma+md+b)-(ma+b) C(ma+md+b)-(ma-b) D y2=m(a+d+b)

OpenStudy (mera8421):

HELP

OpenStudy (holsteremission):

If \(y_2=ma+md+b\) and \(y_1=ma+b\), then \(y_2-y_1=(ma+md+b)-(ma+b)=md\).